OpenStudy (anonymous):

Find the derivative: \(\ \frac{2}{(1+e^{-x})} \)

OpenStudy (anonymous):

what's the derivative of e^x

OpenStudy (anonymous):

e^x

OpenStudy (anonymous):

quotient rule if u want.

OpenStudy (anonymous):

What's confusing me is with that wouldn't that result in multiplication by 0?

OpenStudy (anonymous):

Or is that okay?

OpenStudy (anonymous):

try it out, let me know what u get.

hartnn (hartnn):

or you could use chain rule, given the formula d/dx(1/x) = -1/x^2

OpenStudy (anonymous):

d/dx(2/(1+e^(-x))) = (2 e^x)/(e^x+1)^2

OpenStudy (anonymous):

how did you do that? the drawing

OpenStudy (anonymous):

screenshot.whatever.